OpenStudy (anonymous):

There are 238 juniors at high school. The ratio of girls to boys in the junior class is 3:4. How many juniors are girls? How many are boys?

OpenStudy (anonymous):

3x + 4x = 238 7x = 238 x = 34 3 * 34 = 102 girls 4 * 34 = 136 boys

OpenStudy (anonymous):

wow i knew that... thanks again